# Enhancing Sequential Recommendations through Multi-Perspective Reflections and Iteration

### v1

@@ Abstract.
Sequence recommendation (SeqRec) is a typical task aimed at predicting the next interaction item for users, focusing on understanding user interactive intents and collaborative filtering information. Currently, large language models (LLMs) show great potential in recommendation scenarios through prompt-based, fixed reflection library, and fine-tuning approaches. However, these approaches face challenges such as lack of supervision, inability to optimize reflection sources, inflexibility to diverse users, and high computational costs.
Despite achieving encouraging performance, current studies solely exploit reflections on users’ explicit preferences (e.g., item titles). They ignore reflecting on implicit preferences and collaborative filtering information, which prevents them from capturing the preference shifts embedded in these signals, neglecting reflections on other aspects of preferences(such as implicit preferences (e.g., brands) and collaborative filtering information. As a result, it struggles to capture users’ dynamic preferences. Moreover, current methods lack reflection evaluation and iteration, which inevitably affects the quality of reflections and leads to sub-optimal recommendation. To address these issues, we propose a Mixture of REflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in sequential recommendations. Specifically, we introduce three reflectors for each user to generate LLMs’ reflections on explicit user preferences, implicit user preferences, and collaborative signals. Within each reflector, we devise a self-improving strategy called refining-and-iteration to evaluate and update the reflections. On top of these reflectors, we introduce a meta-reflector responsible for selecting the most appropriate expert and the corresponding reflections for each user’s current recommendation by applying a contextual bandit algorithm. This ensures that we can capture dynamic user preferences effectively.
Extensive experiments on three real-world datasets demonstrate that our method consistently outperforms state-of-the-art approaches. Additionally, we verify that MoRE requires the least training time and GPU memory compared to all other training LLM methods in SeqRec.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Sequential recommendation (SeqRec), which predicts the next item of interest based on a user’s historical interaction sequence, is crucial for recommender systems. The key to this task is to accurately capture users’ dynamic preferences.
@@
Large language models (LLMs) are emerging as promising recommenders (LLMREC) due to their vast world knowledge and excellent reasoning abilities. Two typical approaches in this context are prompt-based methods and fine-tuning methods. Prompt-based methods (Dai et al., 2023; Hou et al., 2023) pre-construct fixed prompts and exploit in-context learning (Dong et al., 2022) to guide the LLM in reasoning the desired item. Fine-tuning methods (Zheng et al., 2024; Zhang et al., 2024) inject domain knowledge by fine-tuning an LLM on substantial annotated recommendation data. Although these methods achieve encouraging performance, the former suffers from an inability to optimize the prompts based on user feedback, while the latter requires substantial computational resources.
Recently, reflection-based methods (Yao et al., 2023; Wang et al., 2024) have emerged as an alternative approach. As shown in Fig. 1, these methods allow LLMs to reflect on their own outputs based on user history and in turn improve future recommendations. Despite their effectiveness, they suffer from the following challenges in modeling and learning user preferences. First, these methods typically rely on explicit preferences, such as titles, to represent items and expect that LLMs can capture user interest patterns based on these features. However, user interests are often implicitly embedded in the attributes of items. For instance, consider a user who is an avid fan of Apple products. If she purchases an Apple smartphone and an Apple t-shirt, the connection between these items is based on the brand rather than the item titles. The implicit preference, in this case, is the brand “Apple,” which signifies a deeper user preference that is not captured solely by examining the titles of the items. Additionally, collaborative filtering (CF) information, which analyzes user behaviors and preferences to find patterns and similarities between users, is not effectively utilized in current approaches. This omission leads to a loss of valuable insights that can be gained from user interactions and preferences. Second, users’ interests, hidden in their behaviors, are intrinsically dynamic and evolve over time. For example, the recommendation at a given time step may depend on reasoning based on explicit preferences, whereas the recommendation at the next step may need to focus on implicit preferences. Therefore, it is necessary to dynamically model and select explicit preferences, implicit preferences, and CF information to ensure that the recommender system can adapt to the changing nature of user preferences and provide more accurate and relevant recommendations. Third, an ideal reflection update strategy should balance both effectiveness and efficiency, but current methods only manage to achieve one of these qualities. One approach (Wang et al., 2024) pre-establishes a fixed pool of reflection candidates to choose from, which may lack the adaptability required for changing user preferences. Another approach (Yao et al., 2023) updates the reflection candidate set using large amounts of data. While this method is potentially more effective in capturing dynamic interests, it is resource-intensive and less efficient.
@@
Figure 1. Workflow of reflection-based methods for SeqRec. LLM reflections are generated offline based on the user’s interaction history ①. These reflections are then collected ② and used as prompts for future online recommendations ③.
Facing the above challenges, we propose a Mixture of REflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in SeqRec. Specifically, we introduce three reflectors for each user to collect LLMs’ reflections based on user interaction history, each focusing on different aspects: explicit user preferences, implicit user preferences, and collaborative signals. To dynamically update the generated reflections, we devise a meta-reflector with a self-improving strategy called refining and iteration. In the refining step, we evaluate the quality of the reflections by measuring the relative improvement in LLMREC, comparing scenarios with and without reflection prompts. When the improvement exceeds a certain threshold, the reflection is considered effective. In the iteration step, we feed the selected effective reflections into LLMs to generate improved reflections, which can then be sent back to the refining step for a new round of evaluation. As such, this offline, self-improving property ensures the effectiveness and efficiency of the reflection update. Moreover, the meta-reflector operates on top of three reflectors, responsible for selecting the most appropriate reflection for each user’s current recommendation. The selection process is formulated as a contextual bandit problem (Li et al., 2010), which can be efficiently optimized. Experimental results demonstrate that MoRE consistently outperforms traditional recommendation methods and LLMREC methods including prompt-based, fine-tuning-based, and reflection-based with minimal GPU memory usage and training time overhead. In summary, our contributions are as follows:
  • We propose an effective and efficient framework, MoRE, for modeling and learning dynamic user preferences in LLM-based SeqRec.
@@
  • MoRE incorporates three reflectors, each collecting reflections from explicit user preferences, implicit user preferences, and CF signals. It also integrates a meta-reflector, which updates the reflections with a refining-and-iteration strategy and selects the appropriate reflection for LLMREC using a contextual bandit algorithm.
@@ 2.1. Sequential Recommendation
SeqRec aims to predict a user’s next item of interest by modeling their past interactions chronologically. The core challenge lies in accurately capturing users’ dynamic preferences. FPMC (Rendle et al., 2010) is an early attempt that addressed this by combining matrix factorization with a Markov chain to capture user preferences. In recent years, deep learning methods have significantly advanced recommendation performance. For instance, Hidasi et al. (2015); Tang and Wang (2018); Li et al. (2017) employ convolutional and recurrent neural networks to model local and global interaction patterns. Meanwhile, Li et al. (2022, 2023); Zhou et al. (2022) develop sequential recommendation models based solely on multi-layer perceptrons, demonstrating comparable performance. Additionally, Transformer-based models proposed by Kang and McAuley (2018); Sun et al. (2019); He et al. (2021, 2022) offer precise item-to-item relevance modeling through self-attention, achieving state-of-the-art results.
Despite their success, these methods often struggle to fully capture the complexity and fluidity of user preferences. This limitation is due to the lack of world knowledge and comprehensive reasoning ability. Consequently, methods based on large language models have emerged as a promising option for sequential recommendation.
2.2. Large Language Models for Recommendation
@@
Current approaches to adapting LLMs for SeqRec predominantly fall into two categories: prompt-based methods and fine-tuning methods. Prompt-based methods involve manually constructing fixed prompts to guide LLMs in reasoning toward the desired items. For example, LLM4RS (Dai et al., 2023) enhances LLM’s recommendation capabilities by customizing prompts to align with traditional ranking, while LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendations through specialized prompting strategies. However, these fixed prompts rely heavily on human expertise and struggle to adapt to diverse users, often resulting in sub-optimal performance (Dong et al., 2022; Wang et al., 2024). Fine-tuning methods (Geng et al., 2022), on the other hand, formulate SeqRec as a question answering task and fine-tune accordingly. To incorporate recommendation knowledge, they either add special tokens (e.g., Llara (Liao et al., 2024) and LC-Rec (Zheng et al., 2024)) or insert embeddings as texts into the prompt (e.g., BinLLM (Zhang et al., 2024)). Despite these efforts, a significant gap remains between LLM’s pre-training and recommendation-oriented fine-tuning, which necessitates a substantial amount of tuning data to achieve proper alignment.
@@
Recently, Wang et al. (2024) propose Re2LLM, a reflection-based approach that leverages LLMs to reflect on interaction history to enhance future recommendations. However, this approach maintains a static prompt pool and solely reflects on explicit user preferences.
@@ 3.1. Problem Formulation
In SeqRec, let $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ denote the item set, $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ denote the user set, and $\mathcal{S}_{u}=[v_{1}^{u},v_{2}^{u},...,v_{t}^{u}]$ denote the interaction sequence in chronological order for user $u\in\mathcal{U}$, where $v_{t}^{u}\in\mathcal{V}$ is the item that $u$ has interacted with at time step $t$. The goal of this task is to predict the item that $u$ will interact with at $t+1$ given $\mathcal{S}_{u}$:
(1) $\hat{v}=\operatorname*{arg\,max}_{v\in\mathcal{V}}\Pr(v_{t+1}^{u}=v\mid\mathcal{S}_{u}).$
In this study, we focus on the LLM-based recommendation. Following (Dai et al., 2023; Hou et al., 2024), we adopt LLM to directly make the prediction as follows:
(2) $\hat{\mathcal{O}}_{u}=\mathrm{LLM}_{\mathrm{REC}}(P(u,\mathcal{S}_{u},\mathcal{C}_{u})),\hat{\mathcal{O}}_{u}\subseteq\mathcal{C}_{u},$
where $P$ denotes the manually crafted recommendation prompt; $\mathcal{C}_{u}$ and $\mathcal{O}_{u}$ respectively denote the candidate and predicted item list for $u$; all the parameters are concatenated in the text form; and $\mathrm{LLM}_{\mathrm{REC}}$, the LLM for recommendation, is frozen. Please note that reflections are essentially a form of prompts generated by the LLM based on interaction history. Therefore, replacing $P()$ with $Ref()$ in Eq. 2 denotes the reflection-based recommendation.
@@ 3.2. Overall Framework
Figure 2. The overall architecture of our MoRE framework. MoRE incorporates multi-perspective reflectors to generate reflections offline ① and a meta-reflector to self-improve these reflections ② and learn to select the most appropriate one ③. During online recommendation ④, the meta-reflector selects the suitable reflections for the current user.
Multi-Perspective Reflectors. As discussed in Sec. 1, user preferences are often implicitly embedded in the attributes of items and can be inferred through collaborative filtering information. Therefore, MoRE incorporates this module to drive LLMs to generate reflections from the perspectives of a user’s explicit preferences, implicit preferences, and collaborative filtering information. These derived reflections based on historical interactions are considered as hints to enhance LLMs’ future recommendations. The details of this module will be introduced in Sec. 3.3.
@@
Meta-Reflector. This module has two main objectives. First, it maintains three memory banks for each user to collect reflections from the corresponding three reflectors and uses a refining-and-iteration strategy to self-improve the reflections and update the current reflection memories for each user. Second, it is responsible for deciding which of the three perspectives of reflection should be chosen to enhance the current recommendation at each step for the user. This decision-making process is formulated as a contextual bandit problem (Li et al., 2010). These two parts will be elaborated in Sec. 3.4 and Sec. 3.5, respectively.
3.3. Multi-Perspective Reflections Generation

### v2

# MoRE: A Mixture of Reflectors Framework for Large Language Model-Based Sequential Recommendation

@@ Abstract.
Large language models (LLMs) have emerged as a cutting-edge approach in sequential recommendation, leveraging historical interactions to model dynamic user preferences. Current methods mainly focus on learning processed recommendation data in the form of sequence-to-sequence text. While effective, they exhibit three key limitations: 1) failing to decouple intra-user explicit features (e.g., product titles) from implicit behavioral patterns (e.g., brand loyalty) within interaction histories; 2) underutilizing cross-user collaborative filtering (CF) signals; and 3) relying on inefficient reflection update strategies. To address this, We propose MoRE (Mixture of REflectors), which introduces three perspective-aware offline reflection processes to address these gaps. This decomposition directly resolves Challenges 1 (explicit/implicit ambiguity) and 2 (CF underutilization). Furthermore, MoRE’s meta-reflector employs a self-improving strategy and a dynamic selection mechanism (Challenge 3) to adapt to evolving user preferences. First, two intra-user reflectors decouple explicit and implicit patterns from a user’s interaction sequence, mimicking traditional recommender systems’ ability to distinguish surface-level and latent preferences. A third cross-user reflector captures CF signals by analyzing user similarity patterns from multiple users’ interactions. To optimize reflection quality, MoRE’s meta-reflector employs a offline self-improving strategy that evaluates reflection impacts through comparisons of presence/absence and iterative refinement of old/new versions, with a online contextual bandit mechanism dynamically selecting the optimal perspective for recommendation for each user. Experiments on three benchmarks show MoRE outperforms both traditional recommenders and LLM-based methods with minimal computational overhead, validating its effectiveness in bridging LLMs’ semantic understanding with multidimensional recommendation principles. Code: https://github.com/E-qin/MoRE-Rec.
Large Language Model Reflection, Sequential Recommendation
journalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Ithaca, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Language modelsccs: Information systems Personalization
@@ 1. Introduction
Sequential recommendation (SeqRec), which predicts the next item of interest based on a user’s historical interaction sequence, is crucial for recommender systems. The key to this task is to capture users’ dynamic preferences accurately. Large language models (LLMs) are emerging as promising recommenders (LLMREC) due to their vast world knowledge and excellent reasoning abilities. Two typical approaches in this context are prompt-based methods and fine-tuning methods. Prompt-based methods (Dai et al., 2023; Hou et al., 2023) pre-construct fixed prompts and exploit in-context learning (Dong et al., 2022) to guide the LLM in reasoning the desired item. Fine-tuning methods (Zheng et al., 2024; Zhang et al., 2024a; Liao et al., 2024) inject domain knowledge by fine-tuning an LLM on substantial annotated recommendation data. Although these methods achieve encouraging performance, the former suffers from an inability to optimize the prompts based on user feedback, while the latter requires substantial computational resources.
@@
Figure 1. Workflow of reflection-based methods for SeqRec. LLM reflections are generated offline based on interaction and recommendation history. These reflections are then collected as hints to improve online recommendations.
Recently, (Yao et al., 2023; Wang et al., 2024) propose reflection-based methods where LLMs refine recommendations by analyzing prediction against user interaction (Fig. 1). While effective, they face three key challenges:
(1) Ambiguous separation of explicit features and implicit patterns within one single user’s interaction history. Effectively modeling user preferences in SeqRec requires distinguishing between explicit and implicit signals within user interaction histories. Explicit features are reflected in directly observable item features, such as titles, while implicit patterns are inferred from latent factors in user behavior, such as brand loyalty. Current reflection-based recommendation methods (Yao et al., 2023; Wang et al., 2024) predominantly rely on explicit item features to model user interests, assuming that LLMs can effectively capture preference patterns from surface-level semantics alone. However, this reliance lacks analysis or reflection on the deeper, implicit connections that often drive user behavior. As illustrated in Fig. 2, consider a user who has interacted with several electronic products and later purchases a T-shirt. If the recommender focuses solely on explicit item titles, it struggles to analyze the subtle connection between these interactions and fails to adapt to evolving user interests. In this case, the underlying implicit preference is the brand “Apple”, reflecting a deeper pattern that explicit features alone cannot reveal.
(2) Underutilized collaborative filtering (CF) signals across multiple user interaction histories. CF leverages behavioral similarities across users to detect shared interests (e.g., Apple product fans). Current approaches (Yao et al., 2023; Wang et al., 2024) neglect such signals, limiting their ability to exploit collective behavioral patterns for improved recommendations.
(3) Suboptimal reflection update strategies. Existing reflection-based methods (Wang et al., 2024; Yao et al., 2023) mainly maintain a fixed reflection pool, lacking personalized dynamic maintenance based on user interactions. This inflexibility prevents them from adapting to users’ dynamic preferences.
Facing these challenges, we propose MoRE (Mixture of REflectors) to address the limitations through three coordinated innovations. First, on the offline side, we design specialized reflectors to model both intra-user patterns and cross-user patterns: 1) Two intra-user perspectives disentangle explicit features (e.g., product titles) and implicit behavioral patterns (e.g., brand loyalty) from a single user’s interaction history; 2) A cross-user perspective captures collaborative filtering (CF) signals through item-centric analysis of user-item-user similarity relationships. This decomposition directly resolves Challenges 1 (explicit/implicit ambiguity) and 2 (CF underutilization). Second, also on the offline side, our self-improving meta-reflector implements a lightweight two-phase update strategy: evaluating reflection quality via LLMREC performance gains (both presence/absence and old/new reflections), then iteratively generating improved reflections using top-performing candidates in an offline loop. This approach overcomes Challenge 3 (suboptimal updates) while maintaining computational efficiency. Third, on the online side, the meta-reflector dynamically selects optimal reflections per user through contextual bandit optimization (Li et al., 2010). Experiments demonstrate MoRE’s superiority over traditional methods and LLMREC variants (prompt-based, fine-tuning-based, and reflection-based baselines) while maintaining low GPU memory and time costs.
In summary, our contributions are as follows:
  • We first propose a dynamic LLM reflection framework is proposed for sequential recommendation to model and learn dynamic user preferences.
@@
  • MoRE incorporates three reflectors, each collecting reflections from explicit features, implicit patterns, and CF signals. It also integrates a meta-reflector, which updates the reflections with a refining-and-iteration strategy and selects the appropriate reflection for LLMREC using a contextual bandit algorithm.
@@ 2.1. Sequential Recommendation
Figure 2. Examples of LLM reflections from explicit, implicit, and CF perspectives based on user interactions.
SeqRec predicts users’ next interests by modeling historical interactions chronologically, with the key challenge in capturing dynamic preferences (Shen et al., 2023, 2024; Zhang et al., 2024d, b, 2025; Shi et al., 2024; Zhang et al., 2024c; Tang et al., 2025). Recent advances (Hidasi et al., 2015; Tang and Wang, 2018; Li et al., 2017) employ convolutional and recurrent neural networks, while Li et al. (2022, 2023); Zhou et al. (2022) demonstrate MLP-based approaches achieve competitive performance. Kang and McAuley (2018); Sun et al. (2019); He et al. (2021, 2022) utilize transformer-based models for item relevance modeling, achieving SOTA results. Despite progress, existing methods remain limited in holistic preference modeling due to insufficient world knowledge and reasoning capabilities. This has spurred recent exploration of LLM-based approaches for sequential recommendation.
Figure 3. The overall architecture of MoRE framework. MoRE incorporates multi-perspective reflectors to generate reflections offline ① and a meta-reflector to self-improve these reflections ② and learn to select the most appropriate one ③. During online recommendation ④, the meta-reflector selects suitable reflections for the current user.
2.2. LLM-based Recommendation
Current approaches to adapting LLMs for SeqRec predominantly fall into two categories: prompt-based methods and fine-tuning methods. Prompt-based methods involve manually constructing fixed prompts to guide LLMs in reasoning toward the desired items. For example, LLM4RS (Dai et al., 2023) enhances LLM’s recommendation capabilities by customizing prompts to align with traditional ranking,
@@ 2.2. LLM-based Recommendation
while LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendations through specialized prompting strategies. However, these fixed prompts rely heavily on human expertise and struggle to adapt to diverse users, often resulting in sub-optimal performance (Dong et al., 2022; Wang et al., 2024). Fine-tuning methods (Geng et al., 2022; Cui et al., 2022; Zhang et al., 2023b; Bao et al., 2023; Yue et al., 2023; Liao et al., 2024; Lin et al., 2023; Shi et al., 2025b), on the other hand, formulate SeqRec as a question answering task and fine-tune accordingly. To incorporate recommendation knowledge, they either add special tokens (e.g., Llara (Liao et al., 2024), LC-Rec (Zheng et al., 2024)) or insert embeddings as texts into the prompt (e.g., BinLLM (Zhang et al., 2024a)). Despite these efforts, a significant gap remains between LLM’s pre-training and recommendation-oriented fine-tuning, which necessitates a substantial amount of tuning data to achieve proper alignment. Additionally, (Kim et al., 2024; Zhang et al., 2023a; Zhu et al., 2024; Zheng et al., 2024; Zhang et al., 2024a; Shi et al., 2025a) attempt to integrate collaborative filtering information into LLMs. While effective, these methods lack an interpretation and analysis of the CF information, which prevents the integration of CF information with the powerful understanding and reasoning capabilities of LLMs, leading to potential suboptimal performance.
@@
Recently, Wang et al. (2024) propose Re2LLM, a reflection-based approach that leverages LLMs to reflect on interaction history to enhance future recommendations. However, this approach maintains a static prompt pool without decoupled modeling of explicit and implicit preferences and incorporate CF signals.
@@ 3.1. Problem Formulation
In SeqRec, let $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ denotes the item set, $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ denotes the user set, and $\mathcal{S}_{u}^{t}=[v_{u}^{1},v_{u}^{2},...,v_{u}^{t}]$ denotes the interaction sequence in chronological order for user $u\in\mathcal{U}$ up to time step $t$, where $v_{u}^{t}\in\mathcal{V}$ is the item that $u$ has interacted with at time step $t$. The goal of this task is to predict the item that $u$ will interact with at $t+1$ given $\mathcal{S}_{u}^{t}$:
(1) $\hat{v}=\operatorname*{arg\,max}_{v\in\mathcal{V}}\Pr(v_{u}^{t+1}=v\mid\mathcal{S}_{u}^{t}).$
In this study, we focus on the LLM-based recommendation. Following (Dai et al., 2023; Hou et al., 2024), we adopt LLM to directly make the prediction in a ranking fashion as follows:
(2) $\hat{\mathcal{O}}_{u}^{t+1}=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t},\mathcal{C}_{u}^{t+1})),\hat{\mathcal{O}}_{u}^{t+1}\subseteq\mathcal{C}_{u}^{t+1},$
where $\mathcal{C}_{u}^{t+1}$ and $\hat{\mathcal{O}}_{u}^{t+1}$ respectively denote the candidate and predicted item list for $u$ at step $t+1$; all the parameters are concatenated in the text form; $\mathrm{LLM}_{\mathrm{REC}}$, the LLM for recommendation, is frozen; and $P_{\mathrm{REC}}$ denotes the manually crafted recommendation prompt.
More specifically, as illustrated in Fig. 1, we focus on the two-stage reflection-based LLMREC.
In the offline reflection generation stage, given a user $u$, the interaction sequence $\mathcal{S}_{u}^{t-1}$ up to time step $t-1$ and the candidate list $\mathcal{C}_{u}^{t}$, we utilize LLM to make ranking prediction $\hat{\mathcal{O}}_{u}^{t}$ at step $t$:
(3) $\hat{\mathcal{O}}_{u}^{t}=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t}),\hat{\mathcal{O}}_{u}^{t}\subseteq\mathcal{C}_{u}^{t}.$
The LLM then reflects on the predicted list $\hat{\mathcal{O}}_{u}^{t}$ and the ground-truth item $v_{u}^{t}$ to infer $u$’s preferences:
(4) $Ref_{i,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{i}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
where $\mathrm{LLM}_{\mathrm{REF}}$ denotes a frozen LLM responsible for generating reflections; the subscript $i\in\{EP,IP,CF\}$ for prompt $P_{i}$ and reflection $Ref_{i,u}$ corresponds to the “Explicit Preference”, “Implicit Preference”, and “Collaborative Filtering”, which will be detailed in the next section.
In the online recommendation stage, the derived $Ref_{i,u}$ is incorporated into Eq. 2 to enhance future recommendations:
(5) $\hat{\mathcal{O}}_{u}^{t+1}=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t},\mathcal{C}_{u}^{t+1},Ref_{i,u})),\ \hat{\mathcal{O}}_{u}^{t+1}\subseteq\mathcal{C}_{u}^{t+1}.$
The details of recommendation prompt $P_{{REC}}$ are as follows:
Recommendation prompt $P_{REC}$ You are a recommender to recommend items for a specific user. The user interacted with items in the following order: $\langle S_{u}\rangle$. Reflections on the past recommendation attempt for this user (if any): $\langle Ref_{i,u}\rangle$. There are now $|C_{u}|$ candidate items: $\langle C_{u}\rangle$. Please consider the user’s historical interaction sequence (and reflections), select appropriate items from the candidates, and rank them to recommend to the user. Think step by step. Your recommendations:
@@ 3.2. Overall Framework
Multi-Perspective Reflectors (offline reflection). Motivated by the need to decouple explicit/implicit sequential features from user interaction histories and integrate user-item-user collaborative patterns, MoRE employs three perspective-aware reflection processes:
(1) Explicit/Implicit Decoupling: Two distinct reflections analyze a single user’s history to disentangle explicit preferences (e.g., item titles and descriptions) and implicit preferences (e.g., attribute-level sequential patterns).
(2) CF-Aware Integration: A third reflection captures cross-user collaborative signals (e.g., rating trends and user similarity) through item-centric analysis.
As discussed in Sec. 1, these derived reflections based on historical interactions are hints to enhance LLM’s future recommendations. The details of this module will be introduced in Sec. 3.3.
@@
Meta-Reflector (offline iteration and online selection). This module has two main objectives. First, it maintains three memory banks for each user to collect reflections from the corresponding three reflectors and uses a refining-and-iteration strategy to self-improve the reflections and update the current reflection memories for each user. Second, it is responsible for deciding which of the three perspectives of reflection should be chosen to enhance the current recommendation at each step for the user. This decision-making process is formulated as a contextual bandit problem (Li et al., 2010). These two parts will be elaborated in Sec. 3.4 and Sec. 3.5, respectively.
3.3. Offline: Multi-Perspective Reflections